# X-ray Emission Signatures of Neutron Star Mergers

### v1

@@ 2.2.1 Hypermassive Neutron Stars
If $M_{\rm rem}\sim(1.2-1.3)\,M_{\rm TOV}$, the remnant can withstand gravitational collapse by rapid differential rotation—the defining characteristic of an HMNS. Differential rotation is not sustained; as the NS spins, magnetic field lines are wound into strong toroidal components, launching Alfvén waves that carry angular momentum outward, reducing the degree of differential rotation. This process acts on a characteristic timescale (Shapiro, 2000):
$t_{A}=\frac{R}{v_{A}}\\ \approx 1\left(\frac{B_{0}}{10^{14}\text{ G}}\right)^{-1}\left(\frac{R}{10\text{ km}}\right)^{-1/2}\left(\frac{M}{2M_{\odot}}\right)^{1/2}\text{ s}.$ (3)
@@
where $v_{A}=B/(4\pi\rho^{2})^{1/2}$ is the Alfvén speed. As magnetic braking reduces the degree of differential rotation, the HMNS loses centrifugal support. If the remnant mass still exceeds the maximum mass for a uniformly rotating NS (the supramassive limit), the remnant will promptly collapse to a black hole on the timescale of $t_{A}$.
@@
It has been suggested that a Type I GRB jet may be launched during or shortly after the HMNS phase (e.g. Rosswog et al., 2003). Most & Quataert (2023) conducted numerical simulations showing that an HMNS can launch relativistic outflows that may explain precursor emission of Type I GRBs, or in some cases the Type I GRB itself.
2.2.2 Supramassive Neutron Stars
@@
TOV mass limit $M_{\text{TOV}}$ $2.25\,M_{\odot}$
@@
If $M_{\rm rem}\sim(1.0-1.2)\,M_{\rm TOV}$, then magnetic braking alone will not induce collapse. Even with no degree of differential rotation, a rapidly and rigidly rotating “supramassive neutron star” (SMNS) can withstand gravitational collapse until magnetic dipole or gravitational radiation removes sufficient rotation energy. The corresponding spindown luminosity is
@@
where $B_{p}$ is the polar magnetic field strength, $R$ is the radius, $\Omega$ is the angular velocity, $I$ is the moment of inertia, and $\epsilon$ is the ellipticity. The first term represents magnetic dipole spindown, and the second term represents GW spindown.
@@
$L_{\rm sd}(t)\sim L_{\rm sd}(0)\bigg{(}1+\frac{t}{t_{\rm sd}}\bigg{)}^{-2}$ (5)
@@
For SMNSs, the collapsing time is usually close to the spindown timescale when the star is significantly spun down. The detailed collapsing time depends on the NS EoS, $\Omega_{0}$ and $B_{p}$. In this paper, we follow the results of Ai et al. (2020) and adopt $M_{\rm TOV}=2.25M_{\odot}$, $M_{\rm SMNS}=1.2\,M_{\rm TOV}=2.7M_{\odot}$ as an example. For an SMNS with $B_{p}\sim 10^{15}\,\rm G$, the expected collapse time is $t_{c}\sim t_{\rm sd}\sim 300\,\rm s$.
2.2.3 Stable Neutron Star
@@
If $M_{\rm rem}<M_{\rm TOV}$, the remnant will form a stable NS that can survive indefinitely. Shortly after merger, it will experience a brief phase of differential rotation before settling into uniform rotation. The stable NS is governed by the same physical processes and emission mechanisms as an SMNS, except it does not rely on rotational energy to withstand gravitational collapse. The remnant may go on to power additional EM signatures, subject to the spin period and magnetic field strength, which could manifest as sustained emission, similar to radio pulsars, or as episodic bursts, as seen in soft gamma-ray repeaters and anomalous X-ray pulsars.
2.3 Viewing geometry
@@
Figure 2: The geometric configuration of a NS merger system after the merger, which includes a jet zone (red), a free zone (green), and a trapped zone (black). In the jet zone, one sees the brightest region of emission from the jet core. In the free zone, one sees off-axis jet emission and can observe the central engine directly. In the trapped zone, relevant only when considering an opaque ejecta (i.e., a kilonova), one cannot observe the central engine directly but may still see off-axis emissions from the jet and central engine. Adapted from Sun et al. (2017).
@@
In BNS mergers and BH-NS mergers with successful tidal disruption of the neutron star, a fraction of disrupted material can be dynamically ejected. Due to the pre-merger orbital angular momentum, this ejecta is preferentially launched along the equatorial plane, while significantly less material is ejected along the polar (jet) axis. As a result, regardless of the central engine, the observed X-ray emission is strongly influenced by the viewing geometry (see Fig. 2).
@@
If the observer’s line of sight is closely aligned with the jet axis—i.e., the viewing angle $\theta_{v}$ is smaller than the jet opening angle $\theta_{c}$ (the Jet zone)—bright gamma-ray and X-ray prompt emission from the jet can be observed directly. For viewing angles slightly outside the jet ($\theta_{v}>\theta_{c}$) but not obstructed by the dynamical ejecta (the Free zone), fainter prompt gamma-ray/X-ray emission may still be detected, depending on the angular structure of the jet and emission directly from the central engine. For magnetar engines with extended X-ray emission powered by a spindown wind, the emission can be viewed on-axis for both a jet-zone and a free-zone observer. However, if the observer’s line of sight lies within the Trapped zone, where the central engine is obscured by the dense equatorial ejecta, X-ray signals from the central engine would be significantly suppressed or entirely blocked.
@@
$\approx 6\times 10^{7}\left(\frac{\kappa}{1\text{ cm}^{2}\text{ g}^{-1}}\right)\left(\frac{M}{10^{-2}M_{\odot}}\right)$
@@
where $\rho$ is the average ejecta density and $\Delta r\sim vt$ is the ejecta shell width. Initially, the ejecta is optical thick, however, as the ejecta expands, the optical depth decreases, and the trapped zone will eventually become transparent once $\tau$ drops below unity. The corresponding transparency time is
@@
$\approx 1.33\times 10^{5}\left(\frac{\kappa}{1\text{ cm}^{2}\text{ g}^{-1}}\right)^{1/2}$
@@
where $\beta\approx 3$ is a parameter delineating the ejecta density profile (Metzger et al., 2010). Central engine activity occurring at $t>t_{\rm thin}$ can be directly observed even along lines of sight within the trapped zone. In contrast, any emission launched earlier than $t_{\rm thin}$ must diffuse through the ejecta, and will be delayed by the diffusion time:
@@
It is important to note that if the central engine is a magnetar, and its spin-down energy exceeds the kinetic energy of the dynamical ejecta, the ejecta can be accelerated. This acceleration reduces the optical depth of the ejecta more rapidly, potentially making $t_{\rm thin}$ much shorter and enabling earlier visibility of central engine emission (see, e.g., the “mergernova” scenario; Yu et al. 2013; Ai et al. 2022, 2025).
3 X-ray Emissions from NS mergers
3.1 Prompt jet emission
For both BNS and NS-BH mergers, a relativistic jet is likely launched along the axis normal to the merger plane. The dissipation of the jet makes GRB prompt emission. The radiation mechanism of prompt emission is still not fully understood; leading candidates include synchrotron radiation from an optically thin region and quasi-thermal, Comptonized emission near the photosphere. Synchrotron self-Compton, external inverse Compton, and hadronic cascades have been also proposed to contribute to prompt emission.
Here, instead of involving uncertainties to model the prompt emission beginning from fundamental principles, we opt for a phenomenological approach. We model the temporal profile of a GRB pulse with the fast-rise-exponential-decay (FRED) function (Norris et al., 2005),
@@ 3.1 Prompt jet emission
where $A$ is the normalization constant, $\tau_{1}$ and $\tau_{2}$ are the rise and decay timescales, and $\lambda=\exp(2\mu)$, where $\mu=\sqrt{\tau_{1}/\tau_{2}}$. Within this framework, the peak time is $t_{\rm peak}=\sqrt{\tau_{1}\tau_{2}}$, the interval measured between two $1/e$ amplitudes is $\Delta t=\tau_{2}\sqrt{(1+4\mu)}$, and the pulse asymmetry is defined by $\tau_{2}/\Delta t$.
We model the GRB spectrum with a Band function (Band et al., 1993),
$N(E)=A\times\left\{\begin{aligned} &E^{\alpha}\exp\left(\frac{E\,(\beta-\alpha)}{E_{\text{b}}}\right),&\text{for }E<E_{\text{b}}\\ &E_{\text{b}}^{\alpha-\beta}\exp(\beta-\alpha)\,E^{\beta},&\text{for }E\geq E_{\text{b}}\end{aligned}\right.$ (11)

### v2

@@ II.2.1 Hypermassive Neutron Stars
If $M_{\rm rem}\sim(1.2-1.3)\,M_{\rm TOV}$, the remnant can withstand gravitational collapse by rapid differential rotation—the defining characteristic of a hypermassive neutron star (HMNS). Differential rotation is not sustainable indefinitely—as the NS spins, magnetic field lines are wound into strong toroidal components, launching Alfvén waves that redistribute angular momentum outward and reduce the degree of differential rotation. This process acts on the Alfvén timescale (Shapiro, 2000),
$t_{A}=\frac{R}{v_{A}}\\ \approx 1\left(\frac{B_{0}}{10^{14}\text{ G}}\right)^{-1}\left(\frac{R}{10\text{ km}}\right)^{-1/2}\left(\frac{M}{2M_{\odot}}\right)^{1/2}\text{ s},$ (6)
@@
where $v_{A}=B/(4\pi\rho^{2})^{1/2}$ is the Alfvén speed. As magnetic braking reduces the degree of differential rotation, the HMNS loses centrifugal support. If the remnant mass still exceeds the maximum mass for a uniformly rotating NS (the supramassive limit), the remnant will promptly collapse to a black hole within the timescale of $t_{A}$.
@@
The idea that a Type I GRB jet may be launched during or shortly after the HMNS phase has been suggested long ago (e.g. Rosswog et al., 2003). Recently, Most and Quataert (2023) conducted numerical simulations showing that an HMNS can launch mildly-relativistic outflows that may explain precursor emission of Type I GRBs, or in some cases the Type I GRB itself.
II.2.2 Supramassive Neutron Stars
@@
TOV mass limit $M_{\text{TOV}}$ $2.15\,M_{\odot}$
@@
If $M_{\rm rem}\sim(1.0-1.2)\,M_{\rm TOV}$, then magnetic braking alone will not induce collapse. A rapidly and rigidly rotating supramassive neutron star (SMNS) can withstand gravitational collapse by rigid rotational energy alone. Once magnetic dipole and gravitational radiation expels sufficient rotational energy, the SMNS will collapse into a BH. The corresponding spindown luminosity is
@@
where $B_{p}$ is the polar magnetic field strength, $R$ is the radius, $\Omega$ is the angular velocity, $I$ is the moment of inertia, and $\epsilon$ is the ellipticity. The first term represents magnetic dipole spindown, and the second term represents GW spindown. Note that the so-called “dipole spindown” term includes the traditional vacuum dipole spindown component and the magnetospheric wind component, which are of the same order but have different inclination-angle dependencies so that the final expression is insensitive to the inclination angle, see e.g. (Goldreich and Julian, 1969; Spitkovsky, 2006).
@@
$L_{\rm sd}(t)\sim L_{\rm sd}(0)\bigg(1+\frac{t}{t_{\rm sd}}\bigg)^{-2},$ (8)
@@
The collapse time of an SMNS can be approximated by its spindown timescale, corresponding to the time at which the star has lost sufficient rotational support. A detailed calculation of the collapse time depends on the NS EoS, $\Omega_{0}$ and $B_{p}$ (see e.g., Beniamini and Lu, 2021). In this paper, we follow the results of Ai et al. (2020) and adopt $M_{\rm TOV}=2.15M_{\odot}$, $M_{\rm SMNS}=1.25\,M_{\rm TOV}\approx 2.7M_{\odot}$ as a typical model. For an model SMNS with parameters listed in Table 2, the expected collapse time is $t_{c}\sim t_{\rm sd}\approx 300\,\rm s$. We note that the model parameters used here are consistent with methods to obtain independent constraints of GW170817/GRB 170817A from the ejecta mass and jet-launching delay timescale (Margalit and Metzger, 2017; Gill et al., 2019).
II.2.3 Stable Neutron Star
@@
If $M_{\rm rem}<M_{\rm TOV}$, the remnant will evolve as a stable NS that can survive indefinitely. Shortly after merger, there may be a brief phase of differential rotation before settling into uniform rotation. Depending on its spin period and magnetic field strength, the remnant may power additional EM signatures—either as sustained emission, akin to radio pulsars, or as episodic bursts, resembling those from soft gamma repeaters and anomalous X-ray pulsars.
II.3 Viewing geometry
@@
Figure 2: The geometric configuration of a post-merger central engine, which includes a jet zone (red), a free zone (green), and a trapped zone (black). In the jet zone, one sees the brightest region of emission from the jet core. In the free zone, one sees off-axis jet emission and can observe the central engine directly. In the trapped zone, relevant only when considering an opaque ejecta (i.e., a kilonova), one cannot observe the central engine directly but may still see off-axis emissions from the jet and central engine. Adapted from Sun et al. (2017).
@@
In BNS mergers and BH-NS mergers with successful tidal disruption of the neutron star, a fraction of disrupted material could be ejected. This tidal dynamical ejecta is preferentially launched along the equatorial plane, reflecting the orbital angular momentum of the system (Kyutoku et al., 2015; Kawaguchi et al., 2015; Lamb and Kobayashi, 2017; Kawaguchi et al., 2016; Foucart, 2020; Sadeh, 2024).
The final jet structure depends on the unknown delay time between the merger and jet launching. If one assumes a short jet-launching delay timescale of $\lesssim 0.1\rm{\,s}$, the jet would quickly penetrate through the ejecta without forming a significant cocoon and evolve into a Gaussian-like jet profile shortly thereafter (Murguia-Berthier et al., 2014; Geng et al., 2019). For longer jet-launching delay timescales, the jet-ejecta interaction would generate a slow, mildly-energetic cocoon that surrounds the central jet, making a two-component jet structure (Gottlieb et al., 2018; Salafia et al., 2020). In general, regardless of the central engine, the observed X-ray emission is strongly influenced by geometry of the system (see Fig. 2).
@@
If the observer’s LoS is closely aligned with the jet axis—i.e., the viewing angle $\theta_{v}$ is smaller than the jet opening angle $\theta_{c}$ (the Jet zone)—bright gamma-ray and X-ray prompt emission from the jet is beamed directly toward the observer. For viewing angles slightly outside the jet ($\theta_{v}>\theta_{c}$) but not significantly obstructed by any dynamical ejecta (the Free zone), off-axis prompt gamma-ray/X-ray emission may still be detected, depending on the angular structure of the jet and emission from the central engine. For magnetar engines with extended X-ray emission powered by a spindown wind, the emission may be directly observed from both the Jet and Free zones. If the observer’s LoS lies within the Trapped zone, where the central engine is obscured by an optically thick ejecta, any X-ray signals from the central engine would be significantly suppressed.
@@
[…]
@@
where $\rho$ is the average ejecta density and $\Delta r\sim vt$ is the ejecta shell width. Initially, the ejecta is optically thick, however, as the ejecta expands, the optical depth decreases, and the Trapped zone eventually becomes transparent to X-rays. The corresponding transparency time is
@@
[…]
@@
where $\beta\approx 3$ is a parameter delineating the ejecta density profile (Metzger et al., 2010). Eq. 11 is based on the assumption that the density profile is uniform.
Central engine activity during $t>t_{\rm thin}$ can be directly observed along any LoS. In contrast, any emission launched earlier than $t_{\rm thin}$ must diffuse through the ejecta, and will be delayed by the diffusion time,
@@
Still, during $t<t_{\rm thin}$, central engine activity can influence the dynamics and evolution of the ejecta and associated X-ray signals. For a magnetar central engine, spindown would inject energy into the ejecta; if the injected energy exceeds the kinetic energy of the dynamical ejecta, the ejecta will be accelerated and significantly reduce $t_{\rm thin}$ (see, e.g., the “mergernova” scenario; Yu et al. 2013; Ai et al. 2022, 2025). Additionally, the magnetar wind would interact with the dynamical ejecta and drive a shock that can radiate in soft X-rays (Siegel and Ciolfi, 2016a, b). In our discussion, the Trapped zone refers to the optically thick regions where a “naked” magnetar wind may not be directly observed.
III X-ray Emissions from NS mergers
III.1 Prompt jet
For both BNS and NS-BH mergers, a relativistic jet may be launched perpendicular to the orbital plane. Energy dissipation within the relativistic jet powers GRB prompt emission. The radiation mechanism of prompt emission is still not fully understood; leading candidates include synchrotron radiation from an optically thin region and quasi-thermal, Comptonized emission near the photosphere. Synchrotron self-Compton, external inverse Compton, and hadronic cascades have been also proposed to contribute to prompt emission (for comprehensive reviews of proposed mechanisms, see, e.g., Mészáros (2006); Kumar and Zhang (2015); Zhang (2018)). Here, instead of involving uncertainties to model GRB prompt emission beginning from fundamental principles, we opt for a phenomenological approach. The framework described in this section and used to produce prompt X-ray light curves and spectra is packaged as the open-source Python package PromptX (Chen et al., 2025).
We model the temporal profile of a GRB pulse with the fast-rise-exponential-decay (FRED) function (Norris et al., 2005; Hakkila et al., 2008),
@@ III.1 Prompt jet
where $A$ is the normalization constant, $\tau_{1}$ and $\tau_{2}$ are the rise and decay timescales, respectively, and $\lambda=\exp(2\mu)$, where $\mu=\sqrt{\tau_{1}/\tau_{2}}$. Within this framework, the peak time is $t_{\rm peak}=\sqrt{\tau_{1}\tau_{2}}$, the interval measured between two $1/e^{3}$ amplitudes is $\Delta t=\sqrt{9+12\mu}$, and the pulse asymmetry is defined by $\Delta t/(3+2\mu)$.
We model the GRB spectrum with a Band function (Band et al., 1993)
$N(E)=A\times\left\{\begin{aligned} &E^{\alpha}\exp\left(\frac{E\,(\beta-\alpha)}{E_{\text{b}}}\right),&\text{for }E<E_{\text{b}},\\ &E_{\text{b}}^{\alpha-\beta}\exp(\beta-\alpha)\,E^{\beta},&\text{for }E\geq E_{\text{b}},\end{aligned}\right$. (14)